OpenStudy (anonymous):

6m+n=32 m-7n=34 ordered pair using substitution

OpenStudy (mertsj):

From the second equation m = 34+7n Replace m in the first equation with that value and solve.

OpenStudy (anonymous):

n = -6m+32 m-7(-6m+32)=34 m+42m-234=32 43m=266 m=266/43 n=-6(266/43)+32

OpenStudy (anonymous):

m=6 and n=-4??

OpenStudy (mertsj):

yes. You are correct.

OpenStudy (anonymous):

Where did I mess up?

OpenStudy (anonymous):

auh!! thank you!

OpenStudy (mertsj):

@mridik 7(32)=224

OpenStudy (anonymous):

Ah darn it! Thanks